# Measuring Compliance of Consent Revocation on the Web

### v2

@@ Abstract
The GDPR requires websites to facilitate the right to revoke consent from Web users. While numerous studies measured compliance of consent with the various consent requirements, no prior work has studied consent revocation on the Web. Therefore, it remains unclear how difficult it is to revoke consent on the websites’ interfaces, nor whether revoked consent is properly stored and communicated behind the user interface.
@@
Our work aims to fill this gap by measuring compliance of consent revocation on the Web on the top-200 websites. We found that 19.87% of websites make it difficult for users to revoke consent throughout different interfaces, 20.5% of websites require more effort than acceptance, and 2.48% do not provide consent revocation at all, thus violating legal requirements for valid consent. 57.5% websites do not delete the cookies after consent revocation enabling continuous illegal processing of users’ data.
@@
Moreover, we analyzed 281 websites implementing the IAB Europe TCF, and found 22 websites that store a positive consent despite user’s revocation. Surprisingly, we found that on 101 websites, third parties that have received consent upon user’s acceptance, are not informed of user’s revocation, leading to the illegal processing of users’ data by such third parties. Our findings emphasise the need for improved legal compliance of consent revocation, and proper, consistent, and uniform implementation of revocation communication and data deletion practices.
1 Introduction
@@
In the past few years, compliance with privacy and data protection laws on the Web has gained a lot of attention, both from the research community and regulators. The EU Data Protection framework, consisting of the General Data Protection Regulation (GDPR) [1] and ePrivacy Directive (ePD) [2] sets requirements for a legally-valid consent when tracking technologies are deployed on a website or app. Such consent is usually implemented via consent banners and must satisfy seven requirements for validity: it must be prior to any data collection, freely given, specific, informed, unambiguous, readable and accessible, and revocable [1, Art. 4(11), 7] [3]. While numerous studies have evaluated compliance of consent banners with the various consent requirements, such as the presence and visibility of the rejection button  [4, 5, 6], one aspect has so far been overlooked by the research community and the regulators: the requirement of revocable consent.
@@
According to the GDPR, users have the right to revoke consent at any time [1, Art. 7(3), Recital 42]; thus, websites must facilitate the exercise of this right [1, Art. 12(2)] by providing an option to revoke consent and delete related data. Websites that do not allow consent revocation are deemed to be processing data illegally, and run the risk of being fined for not complying with the legal requirements [7, 8, 9, 10, 11]. However, it remains unclear whether websites provide users with compliant methods to revoke their consent, whether revoked consent is properly stored by the websites behind the consent interface, and whether it is communicated to and deleted by third-parties that collected users’ data. Our work aims to fill this gap by measuring the compliance of consent revocation on the Web. We address the following questions:
@@
    Are revocation interfaces compliant with EU data protection laws?
@@
To address these questions, we have set up a team of computer scientists and a legal expert to measure compliance of consent revocation on the Web, making the following contributions:
  1. 1.
@@
  2. 2.
    We provide a method to analyse consent revocation interfaces (§4.1), and apply our analysis on the top 200 websites identifying multiple instances of non-compliance with the legal requirements (§5).
  3. 3.
    We study the effect of revoking advertising and analytics (AA) cookies, finding that on the majority of websites revocation does not decrease the number of AA cookies, thus violating the GDPR (§6).
  4. 4.
    We propose a methodology (§4.2) to evaluate revocation on websites that implement IAB Europe Transparency and Consent Framework (TCF) and the most popular CMP, OneTrust. We analyse 281 such websites and evaluate the storage of consent behind the consent banner interface (§7) and the communication of revocation (§8), thereby detecting multiple non-compliant practices.
  5. 5.
@@
Table I presents the most prevalent (more than 10%) potential violations of the GDPR and ePD that we have detected while analysing both consent revocation interface, and consent stored and shared with third parties behind the user interface.
Violation Prevalence
Revocation via a different interface/medium (§5) 19.87% (32/161)
Two or more steps to revoke vs zero steps to accept (§5) 20.5% (33/161)
@@
Third-parties informed of consent acceptance via HTTP requests, but not informed of consent revocation (§8) 74.2% (101/136)
@@
TABLE I: Summary of most prominent potential violations of consent revocation implementation. In brackets we mark the number of websites with violations and the total number of websites analysed for each case.
2 Background and Related Works
@@
In this section, we provide background on different techniques for consent management, and discuss prior works on compliance in the context of consent banners, opt-out and revocation.
@@
Consent Interfaces and Compliance. Consent banners have become a common method for obtaining consent for online tracking. Recent works [3, 12, 13, 14, 15, 16] have examined consent banners following the implementation of the GDPR and the ePD while numerous others [17, 18, 19, 20, 21, 22, 23, 4, 6, 24, 25] explored the impact of privacy regulations on consent design, and how the websites and CMPs might violate these laws, e.g., by using dark patterns. Tools have been proposed to help users handle consent banners automatically [26, 27, 28, 29, 5, 30, 31, 32]. Several works [5, 14, 6, 33] focused on automated analysis of consent banners at scale.
@@
IAB Europe TCF and Compliance. The current “de-facto” standard for the consent banners in the EU is the IAB Europe Transparency and Consent Framework (TCF) [34]. The current TCF v2.2 [35] defines the pre-defined purposes for processing personal data by Consent Management Platforms (CMPs). If also defines the format to store the user’s choice, called TCString, that includes: (a) list of enabled third-party vendors registered within the TCF; (b) list of enabled purposes among the pre-defined purposes [36]. The TCF standard does not specify which of the purposes require user consent. According to GDPR and ePD, only purposes that do not require user consent (and users’ explicit interaction with the CMP) may be enabled in TCString  by default. Matte et al. [37] analysed which purposes in TCF v1.1 and v2.0 require consent according to the GDPR and ePD; however, this analysis was not extended to TCF v2.2, which we perform in this paper.
@@
Once a user has made their choice on the CMP interface, the CMP is required to store the corresponding TCString in the browser (however, the specific storage mechanism is not specified in TCF v2.2), or implement an API for third-parties to check the TCString. The TCString can be accessed via the __tcfapi function by third-parties in first-party context or via the JavaScript postMessage API to communicate with a special __tcfLocator iFrame. Following TCF v1.1 [38], TCString can also be shared in the outgoing HTTP requests via URL-based methods. However, TCF v2.2 does not specify any URL parameters to be used for this.
@@
Prior works [4, 37, 39, 40] have focused on the analysis of IAB Europe TCF and its consent implementation. However, none of these works studied how CMPs manage consent revocation, which is one of the main objectives of our work.
@@
Opt-out and Consent Revocation. While multiple studies detected and analysed the presence of a reject button on consent banner interfaces [4, 6, 5], none of the prior works evaluated the interfaces that allow revocation of a given consent within the EU data protection framework. Nevertheless, since the updates of the CPRA regulation in California, multiple studies have analysed the implementation of the right to opt out of selling or sharing user data within websites. Tran et al. [41] propose a methodology to measure compliance of opt-out links’ wordings automatically found on websites subject to CCPA and CPRA [42]. Liu et al. [43] measure the impact of opting out in the presence of advertisers and tracking mechanisms.
Aziz et al. [25] assessed the IAB CCPA Compliance Framework (analogous to IAB TCF in Europe) and detected that opt-out signals are not honored on websites. A similar latest work by Du et al. [44] detected violation regarding withdrawal interfaces in mobile applications.
@@
Other studies evaluated user perceptions when interacting with opt-out mechanisms. As such, Habib et al. [45] evaluated the usability of data deletion and opt-out options related to email communication and targeted advertising. Habib et al.[46] evaluated how different instructions about revocation in the banner’s text impact users’ ability to find revocation options.
@@
None of these works, however, studied how difficult it is for the user to revoke consent, and whether revocation interfaces comply with the EU laws. Additionally, no prior work has shed light on whether websites correctly store and manage revoked consent, and inform all the third-parties that were previously informed of consent acceptance.
3 Legal Compliance for Consent Revocation
@@
EU Data Protection framework provides legal principles and requirements that can be applied to websites and specifically to consent revocation. The GDPR applies to the processing of personal data [47], and the ePrivacy Directive (ePD) [2] provides supplementary rules, particularly, for the use of tracking technologies [48]. Whenever cookies and other tracking technologies are stored and read from the user’s device, the ePD [2, Art. 5(3)] requires websites to request user’s consent when tracking is used for purposes that are not strictly necessary for the service provided, e.g., advertising [49, 50]. Some purposes are exempted of consent, e.g, functional purposes required for a website to operate [2, Recital 66]. The way to assess with certainty whether consent is required is to analyze the purpose of each tracker on a given website [50, 51].
@@
Websites must not process personal data in a unjustifiably detrimental, discriminatory, unexpected or misleading way [1, Art. 5(1)(a)][52, para 17].
@@
Websites must implement technical and organisational measures, and safeguards efficiently to enable the exercise of the revocation right [53, para 68], [1, Art. 25(1)]; [53, para 70]. Revocation options should be provided in an objective and neutral way, avoiding any deceptive or manipulative language or design [54, 55, para 16]. If a website requires more effort to revoke than to give consent, it is deploying a dark pattern [55, para 30].
@@
Websites must be able to demonstrate that revocation is effective, i.e., performed easily and effectively [1, Arts. 5(2), 24(1), Rec. 74].
4 Methodology
In this section, we describe how the data was collected and the analysis we performed to answer our research questions. We built a semi-automated crawler based on a Selenium-instrumented Chromium v122.0.6261.128. We used the crawler to collect data in April-June, 2024111The code and dataset is available as supplementary material, see 10. within the EU. Table II presents an overview of the three datasets we collected, which are further explained in this section.
Figure 1: Data collection pipeline: to address RQ1, we collect screenshots and label a website as shown in violet boxes; for RQ2 we collect cookies at each stage and also at Rejection stage in a similar way (not shown in this figure).
4.1 Datasets for RQ1 and RQ2
We have implemented the data collection pipeline shown in Figure 1 to capture the elements of the user interface that need to be interacted with in order to exercise consent revocation. At each stage of our analysis, we save background data to further analyse cookies, consent storage and communication. Figure 2 presents the details of the background data collection.
@@ 4.1 Datasets for RQ1 and RQ2
DS1: 105 domains without CMPs RQ1, RQ2
DS2: 56 domains with CMPs RQ1, RQ2,
@@
DS3: 225 domains with CMPs RQ3, RQ4
TABLE II: Datasets used for RQs.
4.1.1 Website selection
We used the top-200 domains from Tranco [64, 65] for analysing the interface of revocation options. 161 domains were reachable while 39 domains did not display a webpage because they were either CDNs, were blocked or returned failures. We detected the CMP by decoding the TCString returned by the __tcfapi following the methodology of Matte et al. [4] after accepting consent. We additionally detect the websites using OneTrust CMP by the presence of OptanonConsent cookie, specific for this CMP. Of the 161 domains, 56 used TCF and/or OneTrust CMP (labeled “with CMP”), which we could detect.
4.1.2 Analyzing interfaces for consent revocation (RQ1)
We collected the data in a semi-automated manner as shown in Figure 1: we save screenshots to record the user interfaces involved in reaching a revocation option. While recent works [66, 67, 68] provide automated tools to select initial consent options on websites (specifically, button and toggle HTML elements) based on user choices, our analysis shows that revoking a previously given consent may require navigating multiple links on the page to reset the options once the initial banner is closed. Moreover, it is not always clear where the consent modification options are located on the website, and whether such options are present. Therefore, revocation options are difficult to find and interact with automatically due to heterogeneity of revocation implementations, as we show in this paper. The data is collected in four stages:

### v3

# Johnny Can’t Revoke Consent Either: Measuring Compliance of Consent Revocation on the Web

@@ Abstract.
The EU General Data Protection Regulation (GDPR) requires websites to facilitate the right to revoke consent from Web users. Prior works have examined consent management by auditing that user choices are correctly stored, and comparing cookies set upon acceptance versus rejection to assess compliance. While these studies measured compliance of consent with respect to the various consent requirements, no prior work has studied consent revocation on the Web. Therefore, it is unclear how difficult it is to revoke consent on the websites’ interfaces, and whether the revoked consent is properly stored and communicated behind the user interface.
@@
Our work aims to fill this gap by measuring compliance of consent revocation on the Web on Tranco’s top-200 websites. We found that 19.87% of websites make it difficult for users to revoke consent throughout different interfaces, 20.5% of websites require more effort than acceptance, and 2.48% do not provide consent revocation at all, thus violating EU legal requirements for valid consent. 57.5% websites do not delete the cookies after consent revocation enabling continuous illegal processing of users’ data.
@@
Further, we analyzed 281 websites implementing the IAB Europe Transparency & Consent Framework, and found 22 websites that store a positive consent despite user’s revocation. Surprisingly, we found that on 101 websites, third parties that have received consent upon user’s acceptance, are not informed of revocation, leading to the illegal processing of users’ data by such third parties according to EU laws. Our findings emphasize the need for improved legal compliance of consent revocation, and proper, consistent, and uniform implementation of revocation communication to third-parties.
journalyear: 2025journalvolume: YYYYjournalnumber: Xdoi: XXXXXXX.XXXXXXX
1. Introduction
@@
In recent years, compliance with privacy and data protection laws on the Web has gained a lot of attention, both from the research community and regulators. The EU Data Protection framework, consisting of the General Data Protection Regulation (GDPR) (European Parliament and Council of the European Union, 2016) and ePrivacy Directive (ePD) (ePD-09, 2009) sets requirements for a legally-valid consent when tracking technologies are deployed on a website or app. Such consent is usually implemented via consent banners and must satisfy seven requirements for validity: it must be prior to any data collection, freely given, specific, informed, unambiguous, readable and accessible, and revocable (European Parliament and Council of the European Union, 2016, Art. 4(11), 7)(Santos et al., 2020, para 17). While numerous studies have evaluated compliance of consent banners with the various consent requirements, such as the presence and visibility of the rejection button  (Matte et al., 2020a; Bollinger et al., 2022; Nouwens et al., 2020), one aspect has so far been overlooked by the research community and the regulators: the requirement of revocable consent, also known as consent withdrawal, which entails that the user has the option to change a prior preference regarding trackers due to the reversible nature of consent decisions (Directorate General Justice, European Commission, 2013).
According to the GDPR, users have the right to revoke consent at any time (European Parliament and Council of the European Union, 2016, Art. 7(3), Rec. 42). Accordingly, websites must facilitate the exercise of this right (European Parliament and Council of the European Union, 2016, Art. 12(2)) by providing an option to revoke consent. Upon revoking consent, websites must subsequently comply with an additional obligation to delete data previously processed on the basis of that consent, and without undue delay even if the user did not explicitly exercise their right to request data deletion (European Parliament and Council of the European Union, 2016, Art. 17(1)(b)). This paper focuses exclusively on the right to revoke consent. Websites that do not allow consent revocation are deemed to be processing data illegally, and run the risk of being fined for not complying with the legal requirements (NOYB, 2024; for Digital Government, 2023; CNIL, 2023; (2019), UODO; DPA, 2024).
@@
However, it remains unclear whether websites provide users with compliant methods to revoke their consent, whether revoked consent is properly stored by the websites behind the consent interface, and whether it is communicated to third-parties that collected users’ data. Our work aims to fill this gap by measuring the compliance of consent revocation on the Web, addressing the research questions:
@@
    Are the revocation interfaces on websites compliant with EU data protection laws?
@@
To address these questions, we have set up a team of several computer scientists and a legal expert to measure compliance of consent revocation on the Web, making the following contributions:
  1. (1)
@@
  2. (2)
    We provide a method to audit compliance of consent revocation interfaces (§4.1), and apply our analysis on Tranco’s top-200 websites identifying multiple instances of non-compliance with the legal requirements (§5).
  3. (3)
    We study the effect of revoking advertising and analytics (AA) cookies, finding that on the majority of websites, revocation does not decrease the number of AA cookies, thus violating the GDPR (§6).
  4. (4)
    We propose a methodology (§4.2) to evaluate revocation on websites that implement IAB Europe Transparency and Consent Framework (TCF) and the most popular Consent Management Platform, OneTrust. We analyse 281 such websites and evaluate the storage of consent behind the consent banner interface (§7) and the communication of revocation (§8), thereby detecting multiple non-compliant practices.
  5. (5)
@@
Violations within the interface and cookies in the top-200 websites Prevalence
Revocation via a different interface/medium (§5) 20.25% (32/158)
Two or more steps to find revocation interface vs zero steps to accept (§5) 20.8% (33/158)
@@
Violations beyond the interface in 281 websites with CMPs Prevalence
@@
Third-parties informed of consent acceptance via HTTP requests, but not informed of revocation (§8) 74.2% (101/136)
@@
Table 1. Summary of most prominent potential violations of consent revocation implementation. In brackets, we mark the number of websites with violations and the total number of websites analysed for each case.
Table 1 summarizes the most prevalent (more than 10%) potential violations of the GDPR and ePD that we detected while analysing both the consent revocation interface, cookies, and the consent storage and sharing with third-parties. Among the top-200 websites, we find that 22.7% (36) websites are compliant with GDPR and ePD within their revocation interfaces and management of cookies (see remaining potential violations in rows 1-3), and the compliance is more common in websites that employ CMPs. We, therefore, further investigate websites with CMPs for possible non-compliance. Of the 281 with CMPs, 251 provided revocation options – among them, we further check if a positive consent is stored even after revocation and if the consent modification is communicated to third-parties that were informed of acceptance. Overall, we find that 52.6% out of 251 websites providing revocation do not have a possible violation beyond the interface, while the remaining websites either had positive consent after revocation or communicated an initial positive consent to third-parties but did not communicate negative consent after revocation (rows 4-6).
2. Background and Related Works
@@
In this section, we provide background on different techniques for consent management and discuss prior works on compliance in the context of consent banners, opt-out, and revocation.
@@
Consent Interfaces and Compliance. Consent banners have become a common method for obtaining consent for online tracking. Recent works (Santos et al., 2020; Bielova et al., 2024; Degeling et al., 2019; Hils et al., 2020; Kampanos and Shahandashti, 2021; Trevisan et al., 2019) have examined consent banners following the implementation of the GDPR and the ePD while numerous others (Gray et al., 2021; Giese and Stabauer, 2022; Ma and Birrell, 2022; Soe et al., 2020; Mathur et al., 2021; Machuletz and Boehme, 2020; Sanchez-Rola et al., 2019; Matte et al., 2020a; Nouwens et al., 2020; Leenes and Kosta, 2015; Aziz and Wilson, 2024) explored the impact of privacy regulations on consent design, and how the websites and CMPs might violate these laws, e.g., by using dark patterns. Tools have been proposed to help users handle consent banners automatically (I dont care about cookies, [n. d.]; Ninja Cookie, [n. d.]; AU, 2022; superagent, [n. d.]; Bollinger et al., 2022; Jesus and Pandit, 2022; Kulyk et al., 2018; Pandit et al., 2019). Several works (Bollinger et al., 2022; Hils et al., 2020; Nouwens et al., 2020; Kirkman et al., 2023) focused on automated analysis of consent banners at scale.
@@
IAB Europe TCF and Compliance. The current “de-facto” standard for the consent banners in the EU is the IAB Europe Transparency and Consent Framework (TCF) (Europe, 2025). The current TCF v2.2 (IAB, 2025b) defines the pre-defined purposes for processing personal data by Consent Management Platforms (CMPs). If also defines the format to store the user’s choice, called TCString,that includes:(a) list of enabled third-party vendors registered within the TCF; (b) list of enabled purposes among the pre-defined purposes (IAB, 2025a).
The TCF standard does not specify which of the purposes require user consent. According to GDPR and ePD, only purposes that do not require user consent (and users’ explicit interaction with the CMP) may be enabled in TCString  by default. Matte et al. (Matte et al., 2020b) analysed which purposes in TCF v1.1 and v2.0 require consent according to the GDPR and ePD; however, this analysis was not extended to TCF v2.2, which we perform in this paper.
@@
Once a user has made their choice on the CMP interface, the CMP is required to store the corresponding TCString in the browser (however, the specific storage mechanism is not specified in TCF v2.2), or implement an API for third-parties to check the TCString. The TCString can be accessed via the __tcfapi function by third-parties in first-party context or via the JavaScript postMessage API to communicate with a special __tcfLocator iFrame. Following TCF v1.1 (Lab and Europe, 2018), TCString can also be shared in the outgoing HTTP requests via URL-based methods. However, TCF v2.2 does not specify any URL parameters to be used for this.
@@
Prior works (Matte et al., 2020a, b; Smith et al., 2024; Zhang et al., 2024) have focused on the analysis of IAB Europe TCF and its consent implementation. However, none of these works studied how CMPs manage consent revocation, which is one of the main objectives of our work.
@@
Consent Rejection and Revocation. While multiple studies (Matte et al., 2020a; Nouwens et al., 2020; Bollinger et al., 2022) detect and analyze the presence of a reject button on consent banners, none of the prior works evaluated the interfaces that allow revocation of a given consent within the EU data protection framework. Nevertheless, since the updates of the CPRA regulation in California, multiple studies have analysed the implementation of the right to opt out of selling or sharing user data within websites. Tran et al. (Tran et al., 2024) propose a methodology to measure compliance of opt-out links’ wordings automatically found on websites subject to CCPA and CPRA (California State Legislature, 2018). Liu et al. (Zengrui Liu and Saxena, 2024) measure the impact of opting out in the presence of advertisers and tracking mechanisms. Aziz et al. (Aziz and Wilson, 2024) assessed the IAB CCPA Compliance Framework (analogous to IAB TCF in Europe) and detected that opt-out signals are not honored on websites. A similar latest work by Du et al. (Du et al., 2024) detected violations regarding withdrawal interfaces in mobile applications.
@@
Other studies evaluated user perceptions when interacting with opt-out mechanisms. As such, Habib et al. (Habib et al., 2019) evaluated the usability of data deletion and opt-out options related to email communication and targeted advertising. Habib et al.(Habib et al., 2022) evaluated how different instructions about revocation in the banner’s text impact users’ ability to find revocation options.
@@
None of these works, however, studied whether the revocation interfaces comply with the EU laws. Additionally, no prior work has shed light on whether websites correctly store and manage revoked consent, and inform all the third-parties that were previously informed of consent acceptance.
3. Legal Compliance for Revocation
@@
The EU Data Protection Framework provides legal principles and requirements that can be applied to websites and specifically to consent revocation. The GDPR applies to the processing of personal data (Board, 2007), and the ePrivacy Directive (ePD) (ePD-09, 2009) provides supplementary rules, particularly, for the use of tracking technologies ((2019), EDPB). Whenever tracking data is stored and read from the user’s device, the ePD (ePD-09, 2009, Art. 5(3)) requires websites to request user’s consent when tracking is used for purposes not strictly necessary for the service provided, e.g., advertising (European Data Protection Board , Article 29 Working Party(2012), EDPB; (2013), EDPB); purposes required for a website to operate are exempted (ePD-09, 2009, Recital 66). The way to assess with certainty whether consent is required is to analyze the purpose of each tracker on a given website ((2013), EDPB; Fouad et al., 2020).
@@
Websites must not process personal data in a unjustifiably detrimental, discriminatory, unexpected or misleading way (European Parliament and Council of the European Union, 2016, Art.5(1)(a))(European Data Protection Board, 2020a, para 17).
@@
Websites must implement organisational measures and safeguards efficiently to enable the exercise of the revocation right (European Data Protection Board, 2020b, para 68), (European Parliament and Council of the European Union, 2016, Art. 25(1)); (European Data Protection Board, 2020b, para 70). Revocation options should be provided in an objective and neutral way, avoiding any deceptive or manipulative language or design (European Data Protection Board, 2024, 2022, para 16). If a website requires more effort to revoke than to give consent, it is deploying a dark pattern (European Data Protection Board, 2022, para 30).
@@
Websites must demonstrate that revocation is performed easily and effectively (European Parliament and Council of the European Union, 2016, Arts. 5(2), 24(1), Rec. 74).
4. Methodology
Next, we describe the data collection and analysis methodology to answer our research questions. We built a semi-automated crawler based on a Selenium-instrumented Chromium v122.0.6261.128. We used the crawler to collect data between March and June 2024, within the EU. Table 2 presents an overview of the three datasets we collected, which are further explained in this section.
Figure 1. Data collection pipeline: to address RQ1, we collect screenshots and label a website as shown in violet boxes; for RQ2 we collect cookies at each stage and also at Rejection stage in a similar way (not shown in this figure).
4.1. Datasets for RQ1 and RQ2
@@
DS1: 105 domains without
CMPs
RQ1, RQ2
DS2: 56 domains with
CMPs
RQ1, RQ2,
@@
DS3: 225 domains with
CMPs
RQ3, RQ4
Table 2. Datasets used for RQs.
We have implemented the data collection pipeline shown in Figure 1 to capture the elements of the user interface that need to be interacted with in order to exercise consent revocation. At each stage of our analysis, we save background data to further analyse cookies, consent storage, and communication. Figure 2 presents the details of the background data collection.
4.1.1. Website selection
We analyzed the revocation interface options on Tranco top-200 domains (Le Pochat et al., 2019; Tranco, 2023). 158 domains were reachable while 42 did not display a webpage because they were either CDNs, blocked or returned failures. We detected CMPs by decoding the TCString returned by __tcfapi following the methodology of Matte et al. (Matte et al., 2020a), after accepting consent. We detect the websites using OneTrust CMP by the presence of OptanonConsent cookie, specific for this CMP. Of the 158 domains, 56 used TCF and/or OneTrust (labeled “with CMP”), which we could detect.
4.1.2. Analyzing interfaces for consent revocation (RQ1)
We collected the data (including cookies, localstorage, request-response pairs and screenshots) in a semi-automated manner as shown in Figure 1.Manual effort was required to browse and locate revocation options, while all technical data (cookies/storage, screenshots, logs) were collected automatically. Automating this would require robust ML/NLP models, which may not generalize well due to the diversity in UIs and revocation paths, as discussed in results of RQ1, section 5.1.1. While we manually visited and navigated the websites, the data was collected automatically in the background. We took screenshots to record the user interfaces involved in reaching a revocation option. While recent works (Khandelwal et al., 2023, 2021; Bouhoula et al., 2024) provide automated tools to select initial consent options on websites (specifically, button and toggle HTML elements) based on user choices, our analysis shows that revoking consent may require navigating multiple links on the page to reset the options once the initial banner is closed. Moreover, the location of the consent modification options on websites are non-standard, and it is not clear whether (and where) such options are present. Therefore, revocation options are difficult to find and interact with automatically due to the heterogeneity of revocation implementations. We collected this data in four stages: